OpenStudy (anonymous):

what is the generic equation for a exopential function? and how would you figure out what i sin or cos function is?

OpenStudy (anonymous):

do you mean e^x written as a power series

OpenStudy (anonymous):

e^x can be written as 1 x + x^2/2! + x^3/3!+.... by McLaaurins theorem Also series for sin x and cos x can be written as series using the same theorem

OpenStudy (anonymous):

sorry - that should start 1 + x + x^2/2!

OpenStudy (anonymous):

thanks, all i understood was the e^x part but thats all i think i need to know. so all it would be is like 2^x or wtvr?

OpenStudy (anonymous):

That i sin, it's that French guy DeMoirve.

OpenStudy (anonymous):

what chagunanas?

OpenStudy (anonymous):

yes 2^x, 3^x and so on are exponential function

OpenStudy (anonymous):

thx a million!!!!!!

OpenStudy (anonymous):

3-4i=\[5(\cos \theta +i \sin \theta)\]=\[5(\cos 307 degrees +i \sin 307 degrees)\]